OpenStudy (anonymous):

Vector Space: Let V(2) be the set of all 2 x 3 matrices. Verify that V(2) is a vector space.

OpenStudy (anonymous):

your job is to check the axioms for a vector space one at a time

OpenStudy (anonymous):

so for example you check that if \(A,B\) are two 2 by 3 matrices, then \(A+B=B+A\) that is, that is commutative. this follows directly from the definition of matrix addition and the fact that real numbers are comx+y = y+x.mutative

OpenStudy (anonymous):

what is the multiplicative identity... how do I check if that exists, just multiply by a 3 x 2 matrix full of 1's???

OpenStudy (anonymous):

it is an additive identity you are looking for, which is the zero matrix

OpenStudy (anonymous):

matrix multiplication is not defined here

OpenStudy (anonymous):

vector space axioms have nothing about multiplying two vectors, only scalar multiplication

OpenStudy (anonymous):

well i found where it lists the properties of a vector space in my book... and it states. 1. addition is commutative 2. addition is associative 3. additive identity 0 exists 4. additive inverse exists 5. multiplication is associative 6. multiplicative idenity exists 7. distributive law for scalars 8. distributive law for vectors

OpenStudy (anonymous):

do i have to prove all 8 just to show that matrix is a vector space

OpenStudy (anonymous):

"multiplication is associative" means scalar multiplication

OpenStudy (anonymous):

you prove that the set of 2 by 3 matrices form a vector space exactly by showing all three axioms hold. but don't get scalar multiplication confused with matrix multiplication you cannot multiply two 2 by 3 matrices

OpenStudy (anonymous):

yeah, i am understanding all the additive properties, and 5 as well, along with 7 and 8, but i do not know how to show property 6.

OpenStudy (anonymous):

again the multiplicative identity means a scalar not a matrix, so it is 1 in this case

OpenStudy (anonymous):

just then its just multiplying 1 by the matrix, not an actual matrix full of ones

OpenStudy (anonymous):

yes. all the properties of multiplication are about scalar multiplication in general you cannot multiply two vectors

OpenStudy (anonymous):

here is a list if you need it http://www.math.ucla.edu/~tao/resource/general/121.1.00s/vector_axioms.html

OpenStudy (anonymous):

okay, yes. thank you, i dont know why it was confusing me... it makes sense, this proof is just ridiculous and long. i thank you for your help... i have one more question i am posting if you are going to be around.

OpenStudy (anonymous):

it is long, but you will admit that all the steps are utterly trivial. they come from the definitions of matrix addition, and the usual properties of real numbers. the exercise is to get you to learn the axioms, not to prove anything difficult

OpenStudy (anonymous):

thank you again. appreciate your help

OpenStudy (anonymous):

yw